On January 2, 2011, Mode Corporation acquired a new machine with an estimated useful life of three years. The cost of the equipment was $40,000 with a residual value of $5,000.
a. Make journal entries of depreciation and relevant ledger accounts under the two depreciation methods listed below. Also show the impact of accumulated depreciation in balance sheet.
In each case, assume that afull year of depreciation was taken in 2011.
1. Straight-line.
2. 200 percent declining-balance.
1
Expert's answer
2020-08-13T12:52:37-0400

1 . straight line method

Depreciation = "\frac{cost value - residual value}{useful life}"

cost value = $40,000 and residual value = $5,000 and useful life = 3 years

Depreciation = "\frac{40,000-5,000}{3}"


Depreciation = "\frac{35,000}{3}"

Depreciation = $11,666.6

this depreciation is for 1 year

and similar for 3 years such as

for 2012

depreciation = $11,666.6

for 2013

depreciation = 11,666.6

BALANCED SHEET





2 . 200% declining balance

first we find depreciation rate :

useful life= 3 , straight line method rate ="\frac{1}{3}"

rate of declining balanced method = ("\frac{1}{3}") 200 = 66.67%

(a) declining balance depreciation for 2011

= 40,000 ("\frac{66.67}{100}")("\frac{12}{12}")

= $26,667

(b)declining balance depreciation for 2012

changing in depreciation rate to 62.50%

=(40,000-26,667)("\frac{62.50}{100}")("\frac{12}{12}" )

= 13,333("\frac{62.50}{100}")

= $8,333

BALANCED SHEET